The width of a rectangle measures 3 kilometers its length is three times its width. What is its perimeter? l
Answer by fractalier(6550)

If the width is 3 km, and the length is three times that, the length must be 9 km.
Perimeter P is found via P = 2L + 2W...here that is
P = 2(9) + 2(3) = 18 + 6 = 24 km